OpenStudy (anonymous):

y varies inversely as twice x. When x = 3, y = 8. Find y when x = 2

OpenStudy (anonymous):

y=12

OpenStudy (anonymous):

y=k/x 8= k/3 k=24 y=24/x y=24/2 y=12

OpenStudy (anonymous):

y=12